OpenStudy (dessyj1):

Calculus 1. Question in the comments

OpenStudy (dessyj1):

\[\int\limits_{none}^{none} \frac{ 11x+4 }{ (3+6x-x ^{2})^{2} } dx\]

OpenStudy (dessyj1):

this is supposed to be an indefinite integral.

OpenStudy (dessyj1):

I cannot easily do u-substitution

OpenStudy (anonymous):

Use partial fraction decomposition \[\frac{ 11x+4 }{(-x^2+6x+3)^2 }=\frac{ Ax+B }{ -x^x+6x+3 }+\frac{ Cx+D }{ (-x^2+6x+3)^2 }\]

OpenStudy (anonymous):

Multiply to eliminate fractions \[11x+4=(Ax+B)(-x^2+6x+3)+Cx+D\]

OpenStudy (dessyj1):

Would you mind showing that to me? I only learned u-substitution.

OpenStudy (dessyj1):

you mean -x^2 instand of x^x right?

OpenStudy (anonymous):

yes I do mean -x². Read up on it here: http://tutorial.math.lamar.edu/Classes/CalcII/PartialFractions.aspx There's a table in the middle of the page that shows what format the decomposed fraction is supposed to do. I'll work through this one.

OpenStudy (anonymous):

Once you have the fractions eliminated, multiply everything out \[11x+4=-Ax^3+6Ax^2+3Ax-Bx^2+6Bx+3B+Cx+D\] Group everything together by the x variable \[11x+4=x^3(-A)+x^2(6A-B)+x(3A+6B+C)+(3B+D)\] *check my multiplication please*

OpenStudy (dessyj1):

checks out.

OpenStudy (anonymous):

Now you can equate coefficients to create a system of equations. On the left, there is no x³ term and on the right the coefficient of x³ is -A, so the equation is \[0=-A\] For x², it's \[0=6A-B\] For x, \[11=3A+6B+C\] For the constant 4 = 3B+D

OpenStudy (dessyj1):

that makes sense

OpenStudy (anonymous):

When I solve the system I get A = 0, B = 0, C = 11, and D = 4, which puts us right where we started :/. PFD is usually a solid method for integrating rational functions. Let me see if I can figure another method

OpenStudy (dessyj1):

Did you solve the system of equations by going off of the fact that you knew A=0?

OpenStudy (anonymous):

yes I did. Ignore this picture for now